# End User Authoring of Personalized Content Classifiers: Comparing Example Labeling, Rule Writing, and LLM Prompting

### v1

@@ 5.2. Supporting Different Content Curation Applications
As discussed, content curation scenarios can range from curating personal feeds to managing community content. Individuals could also have intuitive preferences or gradually develop more well-defined preferences from initial intuitions. Additionally, individuals vary greatly in many dimensions, such as tolerance of undesirable content and the precision-recall trade-off. Future research should therefore only apply our findings after closely examining the unique needs of their contexts. For example, individuals looking to remove undesirable content from their feeds may prefer prompt-based systems over existing rule filters due to their ease of use. In contrast, prompt-based systems might only be desirable for community moderators if they offer comparable transparency to existing rule systems, as moderators are more accountable for their moderation decisions (Song et al., 2023). Additionally, despite their algorithmic similarities, curating desired content requires different systems than removing undesirable content, since users tend to prioritize recall over precision during active exploration and have more fluid and intuitive preferences in this case than in content moderation (Feng et al., 2024). Future work should thus explore how to combine the high recall of writing prompts and the ease of labeling examples into one system for this use case.
In this work, we investigate supporting users to create personal classifiers from scratch but social media users might want to modify existing classifiers. Some platforms have already offered similar but simplified functionalities, allowing users to adjust the sensitivity for predefined concepts like racism and misogyny (Instagram, 2024; Bhargava et al., 2019; Bodyguard, 2024). Compared to a platform-centric definition of these concepts, end users might feel more comfortable customizing classifiers created by friends or family members who share similar interests and moderation preferences (Mahar et al., 2018). Our findings also advocate for more flexible customization options beyond simply adjusting thresholds (Jhaver et al., 2023). For example, we envision users being able to merge labeled data, fork and edit rules, or share prompts.
@@ 7. Conclusion
To communicate feedback to an algorithmic system, end users can primarily leverage three strategies: labeling examples, authoring rules, and writing prompts. Which strategy is the most useful for end users to customize a classifier for content curation? Using personal content moderation as a case study, we conducted a within-subjects experiment with 37 non-programmer participants to compare these three prominent strategies for creating custom classifiers. We found that writing prompts generally allowed participants to create personal classifiers with higher performance more quickly, but not without shortcomings. While prompts could effectively communicate users’ well-defined and general preferences, participants preferred labeling examples to convey their ill-defined but intuitive preferences and authoring rules for their preferences about specific topics or events. Moreover, participants found it challenging to refine their prompts to communicate their nuanced preferences iteratively. Consequently, they often directly added misclassified examples as few-shot examples or attempted to write rule-like prompts. Building on top of our findings, we envision a hybrid approach to custom classifier creation: users could label examples to bootstrap and iterate on their prompt classifiers, while decomposing a complex preference into a rule structure could provide users with greater transparency.
@@ Appendix A Pilot Studies
We conducted pilot studies with our lab members and 5 non-programmers to iteratively implement our three systems. We included non-programmers in these studies to guarantee that our interfaces would be user-friendly and accessible to all potential users.

### v2

@@ 5.2. Supporting Different Content Curation Applications
As discussed, content curation scenarios can range from curating personal feeds to managing community content. Individuals could also have intuitive preferences or gradually develop more well-defined preferences from initial intuitions. Additionally, individuals vary greatly in many dimensions, such as tolerance of undesirable content and the precision-recall trade-off. Future research should therefore only apply our findings after closely examining the unique needs of their contexts. For example, individuals looking to remove undesirable content from their feeds may prefer prompt-based systems due to their ease of use. In contrast, prompt-based systems might be less desirable for community moderators without comparable transparency to existing rule systems, as moderators are more accountable for their moderation decisions (Song et al., 2023). Additionally, despite their algorithmic similarities, curating desired content requires different systems than removing undesirable content, since users tend to prioritize recall over precision during active exploration and have more fluid and intuitive preferences than in content moderation (Feng et al., 2024).
In this work, we investigate supporting users to create personal classifiers from scratch but social media users might want to modify existing classifiers. Some platforms have already offered similar but simplified functionalities, allowing users to adjust the sensitivity for predefined concepts like racism and misogyny (Instagram, 2024; Bhargava et al., 2019; Bodyguard, 2024). Compared to a platform-centric definition of these concepts, end users might feel more comfortable customizing classifiers created by friends or family members who share similar interests and moderation preferences (Mahar et al., 2018). Our findings also advocate for more flexible customization options beyond simply adjusting thresholds (Jhaver et al., 2023). For example, we envision users being able to merge labeled data, fork and edit rules, or share prompts. However, our findings might be less applicable here because users focus less on rapidly initializing a classifier and more on understanding the decision-making boundaries of existing classifiers. In such cases, sharing prompts and rules might be preferred over merging labeled examples, as their high-level nature facilitates quicker understanding.
@@ 7. Conclusion
Existing interactive machine learning systems often make assumptions about end users that do not translate well into the context of content curation. These systems often assume that users are highly motivated to train an algorithm in a single, lengthy session. In contrast, social media users prefer to build their classifiers casually, engaging in shorter, iterative sessions as part of their daily routines. This raises a key question: how can we better support end-users in easily bootstrapping and iterating on classifiers for content curation? To address this, we compared three prominent strategies for creating custom content classifiers: labeling examples for supervised learning, writing and carrying out rules, and prompting a large language model (LLM). From a within-subjects experiment with 37 non-programmer participants, we found that participants preferred different strategies based on their moderation preferences, despite relatively higher performances of prompting LLMs. While prompts could effectively communicate users’ well-defined and general preferences, participants preferred labeling examples to convey their ill-defined but intuitive preferences and authoring rules for their preferences about specific topics or events. Moreover, participants found it challenging to refine their prompts to communicate their nuanced preferences iteratively. Consequently, they often directly added misclassified examples as few-shot examples or attempted to write rule-like prompts. Building on top of our findings, we envision a hybrid approach to custom classifier creation: users could label examples to bootstrap and iterate on their prompt classifiers, while decomposing a complex preference into a rule structure could provide users with greater transparency.
@@ Acknowledgements.
This research was supported by NSF award #2236618. We would like to thank members of the Social Futures Lab at the University of Washington for their invaluable help in this project. We also would like to thank our anonymous reviewers for their insightful feedback. Finally, we would like to express our heartfelt thanks to all the participants who dedicated their time and effort to participate in our study.
@@ Appendix A Pilot Studies
We conducted pilot studies with our lab members and 5 non-technical participants to iteratively implement our three systems. We included non-programmers in these studies to guarantee that our interfaces would be user-friendly and accessible to all potential users.
@@ A.3. Prompt System
Appendix B Examples of User-Created Rules/Prompts
Here we list three examples of prompts that participants authored in our experiments.
  • Remove texts that refer to people as stupid, dumb, idiots, evil, useless, mentally ill, inbreds, nuts, idiot, demented, deranged, scum, retard.
  • Remove off-topic texts such as self-promotion unrelated to the current conversation.
  • Delete texts that disclose private, sensitive personal information (e.g., addresses, phone numbers) unrelated to the discussed topic.
In addition, we list three examples of rules that participants authored in our experiments.
  • Remove comments that include at least one of the following phrases “bitch, stupid, crazy, bitches, stupidity, mad, craziness, dumb, asshole, lunatic, idiot, dick, crackpot, psycho, dumbass, idiots, nutcase, psychos, assholes, moron, freak”.
  • Remove comments that include at least one of the following phrases “republican, Republicans, GOP, conservative, right-wing, Grand Old Party, democrats, conservatives, right-wingers, Republican Party, GOP members, Democratic Party” and at least one of the following phrases “kill, dead, violence”.
  • Remove comments that include at least one of the following phrases “keep them away from African Americans, keep away from blacks, stay away from African Americans, avoid African Americans, keep them from blacks, Republicans are idiots, You bonehead Democrats, Avoid blacks, Stay away from blacks, Keep away from African American, Stay clear of African Americans”.